A fish is 4.3 cm from the front surface of a fish bowl of radius 17 cm. Where does the fish appear to be to someone in air viewing it from in front of the bowl? Do not forget the proper sign.
Answer #1

nw = index of refraction of water = 1.33

na = index of refraction of air = 1

do = object distance = 4.3 cm

di = image distance = ?

R = Radius of bowl = - 17 cm

Using the equation

nw/do + na /di = (nw - na)/R

inserting the values

1.33/4.3 + 1/di = (1.33 - 1)/(- 17)

di = - 3.04 cm